OpenStudy (anonymous):

Donna teaches dance in groups of five students. She rents a hall for the dance classes at $450 per month. She charges $40 per hour per student. Donna teaches four groups of students, and her net earnings per month are $4,350. If each student learns dance for the same number of hours, write an equation that can be used to calculate the number of hours each student learns dance in a month.

OpenStudy (anonymous):

@phi

OpenStudy (sleepyjess):

What do you think you would do first?

OpenStudy (anonymous):

40h * 4 = 4,350 ?? I don't know really...

OpenStudy (anonymous):

I know that there has to be 40h but idk what should come after that

OpenStudy (sleepyjess):

I am thinking it would be something like \(\huge\frac{4350}{40}/4=h\)

OpenStudy (sleepyjess):

do you see where I am coming up with those numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(4350/40) /4 = h write it like this?

OpenStudy (sleepyjess):

That would probably be the best that I could think of but lets see if we might be able to get @dumbcow 's opinion.

OpenStudy (anonymous):

(4350/40) /4 = h (4350/40) = 108.75 108.75 / 4 = 27.1875 h = 27.1875

OpenStudy (sleepyjess):

whoops should be 5 not 4

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

each group (4 groups) has 5 students, so she has like 20 students total

OpenStudy (anonymous):

wait so should the number be 20 in the equation?

OpenStudy (sleepyjess):

oh well then it should be \((\huge\frac{4350}{40})/4*5\)

OpenStudy (sleepyjess):

I must have missed the part about the 4 groups

OpenStudy (anonymous):

(4350/40) /20 = h (4350/40) = 108.75 108.75 / 20 = 5.4375 h = 5.4375

OpenStudy (sleepyjess):

@satellite73 @thomaster

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

just want to make sure that this is correct? :)

OpenStudy (sleepyjess):

@agent0smith

OpenStudy (agent0smith):

h is the number of hours each student does. Four groups, 5 students per group, so 20 students total. 20 students at $40 per hour, so 20*$40 or $800 per hour. 800h would give the $ amount for h, the number of hours Then subtract the money she spends renting the dance hall, $450. 800h - 450 = 4350